OpenStudy (anonymous):

A set of data is normally distributed with a means of 455 and a standard deviation of 25. What percent of data is in the interval 405-455?

OpenStudy (anonymous):

using z scores, do (455-455)/25 and find the p value and do the same for (405-455)/25 then subtract 405 p value from 455 p value. thats the answer

OpenStudy (anonymous):

How do I find the P value?

OpenStudy (anonymous):

455-455/25 = 0. which is 50%. 405-455/25=-2 -2 corresponds to a p value of 0.228. so the percent of data in the interval 405-455 is : .50-.228= .272

OpenStudy (anonymous):

The choices are 47.75% 49.85% 68.3% 34.15%

OpenStudy (anonymous):

P-value is for a different thing. It's for ANOVA, if I recall correctly. This question is somewhere before p-values.

OpenStudy (campbell_st):

405 - 455 is 2 standard deviations below the mean, up to the mean  add the 2 percentages for your answer.